### v1

# in World Models

Abstract
World models aim to understand, remember, and predict dynamic visual environments, yet a unified benchmark for evaluating their fundamental abilities remains lacking. To address this gap, we introduce MIND, the first open-domain closed-loop revisited benchmark for evaluating Memory consIstency and action coNtrol in worlD models. MIND contains 250 high-quality videos at 1080p and 24 FPS, including 100 (first-person) + 100 (third-person) video clips under a shared action space and 25 + 25 clips across varied action spaces covering eight diverse scenes. We design an efficient evaluation framework to measure two core abilities: memory consistency and action control, capturing temporal stability and contextual coherence across viewpoints. footnotetext: Equal contribution. Corresponding author. Furthermore, we design various action spaces, including different character movement speeds and camera rotation angles, to evaluate the action generalization capability across different action spaces under shared scenes. To facilitate future performance benchmarking on MIND, we introduce MIND-World, a novel interactive Video-to-World baseline. Extensive experiments demonstrate the completeness of MIND and reveal key challenges in current world models, including the difficulty of maintaining long-term memory consistency and generalizing across action spaces.
1 Introduction
Table 1: Comparison of World Model Benchmarks. ‘Avg.’ denotes the average number of frames used for memory context and predicted segment in each benchmark. ‘11st-P.’ and ‘33rd-P.’ refer to first person and third person perspectives, respectively. $\emptyset$ denotes the benchmarks without action-based generation (e.g., text–video gen). ‘CharPos.’ refer to the character position. MIND is the first open-domain closed-loop revisited benchmark for evaluating video consistency across both first- and third-person perspectives.
Benchmark CharPos. Fixed Act. (Avg.) Generalized Act. (Avg.) Res./FPS Scenario (image /video )
11st-P. 33rd-P. 11st-P. 33rd-P.
WorldSimBench [32] 1 / - $\emptyset$ $\emptyset$ 1 / - -/- : Minecraft, Driving…
WorldModelBench [22] 1 / - $\emptyset$ $\emptyset$ $\emptyset$ -/- : Humans, Natural…
WorldScore [9] 1 / - $\emptyset$ $\emptyset$ $\emptyset$ -/- : Dining , Passageways…
World-in-World [50] $\emptyset$ $\emptyset$ $\emptyset$ $\emptyset$ 576p/- : Interior environment…
GameWorld [52] 1 / - $\emptyset$ $\emptyset$ $\emptyset$ 720p/- : Minecraft
Lian et al. [26] 65 / 436 $\emptyset$ $\emptyset$ $\emptyset$ 360p/20 , : Minecraft
MIND (Ours) 1.1k/3.4k 1.2k/3.6k 1.3k/3.8k 1.2k/3.7k 1080p/24 , : Landscape, SciFi,
Stylized, Ancient, Urban,
Industrial, Interior, Aquatic
Recent advances in video generation technology have significantly improved the creation of high-fidelity, realistic content, laying a solid foundation for developing sophisticated world models [13, 46, 49, 8, 29]. These models have accelerated advancements across diverse domains, including autonomous driving [25, 30, 44, 21, 51], embodied intelligence [3, 28, 2, 36, 7], and interactive game environments [4, 39, 43, 23, 46], by enabling the generation of complex, diverse, and controllable virtual worlds. Despite these advances, building a reliable world model remains challenging. Beyond visual realism, such models must maintain long-term memory consistency and exhibit accurate action control and robust action generalization across diverse scenarios. Yet, current evaluations mainly focus on visual quality or physical realism, overlooking these essential aspects. Consequently, the field still lacks a comprehensive benchmark to systematically assess memory consistency and action controllability in open-domain environments.
Existing benchmarks primarily focus on evaluating the quality and realism of generated videos, often limited to first-person perspective data collected within a single action space. For instance, WorldScore [9] decomposes scene generation into specific camera motion trajectories to assess video quality, while WorldModelBench [22] evaluates adherence to physical laws to measure world modeling capabilities in application-driven domains. Although Lian et al. [26] introduced a world model memory benchmark, it is limited to Minecraft scenes, lacks open-domain diversity, and depends on loop-based agent data that poorly reflects human behavior. Furthermore, the existing world model benchmark predominantly features first-person perspectives [9, 26], making it challenging to evaluate the ability of world models to simulate motion and poses. In summary, as shown in Table 1, existing benchmarks focus mainly on first-person settings and image-level evaluation, lacking memory consistency assessment and scene diversity. Establishing a comprehensive world model benchmark remains an open and unresolved challenge.
We present MIND, the first closed-loop revisited open-domain benchmark for evaluating memory consistency and action control from both first-person and third-person perspectives across diverse scenarios. MIND focuses on two key abilities of world models: 1) Memory consistency refers to the ability of model to maintain coherent spatial layouts, object identities, and scene attributes over long temporal contexts, ensuring that generated frames remain consistent with past observations. 2) Action control measures how accurately the model executes given control inputs and generalizes these dynamics to new motion ranges or unseen action spaces, reflecting its capacity for precise and adaptable interaction within dynamic environments. Furthermore, the provided videos include frame-level aligned actions, character and camera positions, and image labels, collected from multiple volunteers to capture diverse human behaviors. The dataset contains $250$ high-quality 10801080p, $24$ FPS, frame-level action-aligned videos spanning eight major scene categories, enabling comprehensive evaluation of world models.
To summarize, the contributions of this paper are:
  • Open-Domain Benchmark for World Models. We introduce MIND, the first closed-loop revisited open-domain benchmark at 10801080p / $24$ FPS for evaluating world models from both first-person and third-person perspectives.
  • Evaluation for Memory Consistency and Action Control. We design an efficient framework to assess memory consistency and action control, capturing temporal stability and contextual coherence across viewpoints.
  • Evaluation for Cross-Action Space Generalization. We design various action spaces, including different character movement speeds and camera rotation angles, to evaluate the action generalization capability across different action spaces under shared scenes.
  • The novel Video-to-World baseline, MIND-World. Extensive experiments demonstrate the completeness of MIND and expose key challenges in current world models, such as limited long-term memory consistency and limited generalization across action spaces.
2 Related Work
Figure 1: Overview of the MIND. We build and collect the first open-domain closed-loop revisited benchmark using Unreal Engine 5, supporting both first-person and third-person perspectives with $1080$ p resolution at $24$ FPS.
2.1 Video Generation
Recent advances in video generation models such as SVD [1], Hunyuanvideo [20], Cogvideox [45], Wan [40] and Sora 2 [31] have significantly enhanced video realism, temporal coherence, and controllability, extending generation toward long-horizon, physically plausible scenes. While benchmarks such as VBench [18], VBench-2.0 [54] and EvalCrafter [27] have introduced fine-grained evaluation dimensions (e.g., human fidelity, physics, commonsense).
2.2 World Model
Recent advances in world models have broken down the technical barriers between visual generation and embodied simulation, enabling agents or users to interact in temporally consistent virtual environments. Unlike traditional text-to-video models, world models emphasize long-term memory consistency [42, 16, 41, 12, 48, 24, 53], action-conditioned controlled generation [49, 46, 35, 14, 11] and real-time response [17, 2, 37, 13, 34], evolving into three core research directions. To ensure long-term memory consistency, mainstream existing studies adopt three strategies: pose frame retrieval, context memory compression, and explicit 3D memory representation. Specifically, CAM [48] retrieves context frames based on the field-of-view coverage of pose perspectives, Infinite-World [41] designs a hierarchical pose-free memory compression module to autonomously anchor generated content to distant historical information, and SPMem [42] achieves explicit 3D memory representation by virtue of geometrically anchored long-term spatial memory. For the optimization of action-conditioned controlled generation, GameFactory [49] proposes a multi-stage training strategy integrated with domain adapters, which decouples game style learning from action control to realize scene-generalizable action control and AdaWorld [11] embeds action information into the pre-training process and extracts implicit actions from videos via self-supervision, thus enabling novel action learning under limited conditions. Real-time interaction is a core characteristic of this field, and relevant training paradigms lay a foundation for real-time streaming generation of diffusion-based world models. For instance, Diffusion-Forcing [5] trains diffusion models to denoise token sets with independent per-token noise levels and Self-Forcing [17] performs autoregressive inference with KV caching during training, conditioning the generation of each frame on the model’s own prior outputs. Together, these advances mark a shift from static video synthesis to interactive, temporally consistent world models.
2.3 Evaluation for World Model
The rapid rise of world models has spurred new benchmarks, yet most primarily emphasize scene quality or physical plausibility. WorldScore [9] standardizes camera-trajectory layouts to rate generated video quality. WorldModelBench [22] targets adherence to physical laws in application-driven settings. And WorldSimBench [32] assesses visual realism. However, these efforts under represent two core abilities of world models: long-context memory consistency and action-space generalization across varied controls. In contrast, MIND introduces the first open-domain closed-loop revisited benchmark at 1080p and 24 FPS from both first-person and third-person views, providing unified, efficient protocols to evaluate memory consistency and action control.
3 MIND Benchmark
3.1 Video Source and Environment Settings
To comprehensively evaluate world models across diverse interactive contexts, we construct a large-scale video corpus rendered within Unreal Engine 5. As shown in Figure 5, the benchmark spans $8$ categories covering over $40$ open-domain environments, designed to reflect a wide spectrum of visual and physical dynamics. These include natural (e.g., forest, desert, mountain, ocean), urban (e.g., downtown, residential, industrial), indoor, vehicle, sci-fi, stylized, fantasy, and abstract scenes. As shown in Figure 1 , we construct a systematic data generation pipeline and recruit multiple volunteers to perform both scripted and free-form actions within these environments. We collect $250$ frame-aligned videos at $1080$ p / $24$ FPS. Among them, $200$ videos ($100$ first-person and $100$ third-person, evenly split for training and testing) share the same action space, while the remaining $50$ videos ($25$ per perspective) feature distinct action spaces, providing high-quality and controllable ground truth for evaluation.
3.2 Basic Actions Modeling
In this section, we define a basic action set for modeling both agent translation and camera rotation, which are essential for evaluating action control and scene consistency in world models.
Action Space Definition. We define the action space $\mathcal{A}$ as follows:
$\mathcal{A}=\{W,A,S,D,\uparrow,\downarrow,\leftarrow,\rightarrow\},$
where:
  • $W,A,S,D$ correspond to forward, left, backward, and right movement,
  • $\uparrow,\downarrow,\leftarrow,\rightarrow$ correspond to camera pitch and yaw rotations.
Figure 2: Distribution for Scene Categories and Action Space in MIND. MIND supports open-domain scenarios with diverse and well-balanced action spaces.
Translational Motion. For translational actions, the position of agent $\mathbf{p}_{t}=[x_{t},y_{t},z_{t}]^{\top}$ is updated based on the selected movement direction:
$\mathbf{p}_{t+1}=\mathbf{p}_{t}+\Delta_{p}\cdot\mathbf{v}_{a},$
where $\mathbf{v}_{a}$ is the direction of movement corresponding to the action (e.g., $\mathbf{v}_{W}=[0,0,1]^{\top}$ for forward) and $\Delta_{p}$ is the step size.
Rotational (Camera) Motion. For camera rotation, the orientation $\mathbf{r}_{t}=[\theta_{t},\phi_{t}]^{\top}$ is updated by a small angular increment $\Delta_{r}$:
$\mathbf{r}_{t+1}=\mathbf{r}_{t}+\Delta_{r}\cdot\mathbf{u}_{a},$
where $\mathbf{u}_{a}$ corresponds to the direction of camera rotation (e.g., $\mathbf{u}_{\uparrow}=[0,+1]^{\top}$ for pitch up).
Figure 3: Action Generalization from MIND. Different generalization settings for $\Delta_{p}$ (movement increment) and $\Delta_{r}$ (camera angle increment) are derived from both first-person and third-person perspectives. Each image is captured after the action has been executed for 24 frames.
3.3 Action Space Generalization
In the action modeling framework, the values of $\Delta_{p}$ (for translational motion) and $\Delta_{r}$ (for rotational motion) are not fixed, but can be generalized to accommodate a range of action scales. The action set can be customized to represent different motion scales, as shown in Figure 3. For example, an action with a 0.7-degree rotation and 150 units of translation ($\Delta_{r}=0.7^{\circ},\Delta_{p}=150$) allows for precise control. Larger movements, such as a 1.4-degree rotation with 280 units of translation ($\Delta_{r}=1.4^{\circ},\Delta_{p}=280$), represent broader actions. Conversely, smaller steps like 0.4 degrees of rotation with 100 units of translation ($\Delta_{r}=0.4^{\circ},\Delta_{p}=100$) enable more subtle adjustments, useful for tasks requiring high precision. This flexibility in $\Delta_{p}$ and $\Delta_{r}$ allows the system to adapt to varying levels of control and task requirements.
Action generalization enhances the flexibility of model and realism across diverse scenarios. Thus, world models must adapt to varied action spaces. To assess this adaptability, we collect high-quality, frame-aligned videos from different action spaces within the same scene. Specifically, we configure five combinations of character movement speeds $\Delta_{p}$ and camera rotation speeds $\Delta_{r}$ to generate datasets with diverse action spaces. The setting includes a total of $25$ first-person and $25$ third-person clips, thereby comprehensively and systematically assessing the generalization capability of world models.
3.4 Temporal and Memory Consistency
To evaluate the ability of world models to maintain memory and temporal consistency over time, we introduce a memory revisit strategy, as illustrated in Figure 1. In our setup, a human operator performs predefined actions ($W,A,S,D,\uparrow,\downarrow,\leftarrow,\rightarrow$) within a 3D Unreal Engine 5 environment. The resulting first-person and third-person videos are frame-aligned with action logs and used as ground-truth supervision.
Memory Setup. We define a memory segment as a observed video sequence $\mathcal{M}=\{f_{1},f_{2},\dots,f_{T}\}$, where each frame $f_{t}$ encodes both visual appearance and scene layout. The memory provides contextual grounding that the model must retain when generating subsequent predictions. After observing $\mathcal{M}$, the model receives an action sequence $\mathcal{A}=\{a_{T+1},\dots,a_{T+k}\}$ and is required to predict the future video frames $\hat{\mathcal{V}}=\{\hat{f}_{T+1},\dots,\hat{f}_{T+k}\}$.
Consistency Objective. The model is evaluated on whether the predicted frames $\hat{f}_{T+i}$ remain temporally and spatially consistent with the memorized scene. This includes:
  • Memory Consistency: previously observed objects, layouts, and textures should remain unchanged when revisited through new actions (e.g., returning to the same location should reproduce the same scene appearance);
  • Temporal Consistency: predicted frames should exhibit smooth transitions and coherent dynamics over time, avoiding flickering or sudden structural changes.
Formally, given a revisiting trajectory $\mathcal{A}_{\text{loop}}$ that leads back to a previously seen state, the consistency error can be defined as:
$\mathcal{L}_{\text{mem}}=\|\hat{f}_{t}-f_{t^{\prime}}\|_{2}^{2},$
where $f_{t^{\prime}}$ corresponds to the ground-truth frame at the revisited scene.
3.5 Evaluation
Figure 4: The 10 Symmetric Motion Paths. The blue line represents the original path, and the red line represents the corresponding mirrored path. Each action lasts $24$ frames.
Figure 5: Eight Scene Categories and Action Visualization in MIND. Each category covers multiple representative environments designed to evaluate action-following controllability and history consistency in world models.
Long Context Memory Consistency. Long context memory evaluates the ability of world model to reconstruct previously observed content from contextual memory, reflecting its understanding of scene dynamics and physical laws. Given a full memory sequence $\mathcal{M}=\{f_{1},\dots,f_{T}\}$ and an action sequence $\mathcal{A}=\{a_{T+1},\dots,a_{T+k}\}$, the model generates predicted frames $\hat{\mathcal{V}}=\{\hat{f}_{T+1},\dots,\hat{f}_{T+k}\}$. Ideally, the predicted sequence should match the real sequence $\mathcal{V}=\{f_{T+1},\dots,f_{T+k}\}$ obtained under the same actions. We quantify the long-context memory ability by the mean squared error (MSE) between predicted and ground-truth frames: $\mathcal{L}_{\text{lcm}}=\frac{1}{k}\sum_{i=1}^{k}\|\hat{f}_{T+i}-f_{T+i}\|_{2}^{2}$ where a lower $\mathcal{L}_{\text{lcm}}$ indicates stronger long-term memory retention and reconstruction fidelity.
Generated Scene Consistency. To quantify the world model’s ability to maintain consistency in generated scenes, we introduce a generated scene consistency metric based on $10$ symmetric motion paths (Figure 4), each involving simple translations or rotations lasting $24$ frames. The model moves forward and then retraces the same path in reverse; ideally, frames from the forward (fwd) and reverse (rev) paths should match exactly. We measure this consistency using MSE: $\mathcal{L}_{\text{gsc}}=\frac{1}{k}\sum_{i=1}^{k}\|\hat{f}_{T+i}^{\text{fwd}}-\hat{f}_{T+i}^{\text{rev}}\|_{2}^{2},$ where $\hat{f}_{T+i}^{\text{fwd}}$ and $\hat{f}_{T+i}^{\text{rev}}$ denote the predicted frames from the forward and reverse trajectories, respectively. A lower $\mathcal{L}_{\text{gsc}}$ indicates stronger scene generation consistency and geometric stability.
Action Accuracy. The accuracy of action feedback in world models is central to their precise instruction execution and reliable completion of complex sequential tasks. To evaluate this capability fairly, we unify the predefined action sequences for all models, recover camera trajectories from generated videos via ViPE [15], eliminate scale and coordinate system discrepancies through Sim(3) Umeyama alignment [38], and then calculate translational and rotational relative pose errors. This metric quantifies the accuracy of each model in generating expected frames based on action commands, and is independent of the model’s internal velocity parameters.
Action Space Generalization. World models serve as simulators for domains such as autonomous driving and robotics, where understanding spatial regularities is crucial. We evaluate action space generalization by computing the MSE between generated and ground-truth frames under diverse action settings. Ideally, the model should learn action-space constraints from context and generate videos that follow them with zero-shot consistency.
Visual Quality. We evaluate visual quality from two complementary perspectives: 1) Aesthetic quality. LAION [33] aesthetic prediction model is used to quantitatively evaluate the visual attractiveness of each frame. Trained on large-scale human preference data, it scores composition, color, lighting, realism, and style consistency. Higher scores indicate closer alignment with human aesthetic judgments. 2) Imaging Quality. MUSIQ [19] evaluates perceptual fidelity by detecting artifacts like overexposure, noise, compression, and blur. Trained on the SPAQ [10] dataset, it quantifies image clarity and sharpness as an objective measure of visual quality.
4 Experiment
Figure 6: MIND-World model framework. The parameterized action injection mechanism enables frame-level alignment with input videos and streamlines the process, forming an efficient baseline for Video-to-World training and inference. Furthermore, it allows for unlimited-length inference based on action sequences.
4.1 MIND-World
We present MIND-World, a high-dynamics, real-time, interactive autoregressive video generation pipeline (see Figure 6). Following [13], our training pipeline has three stages: (i) a bidirectional, action-conditioned teacher model, (ii) student initialization from the teacher’s ODE trajectories [47], and (iii) self-forcing DMD distillation [17] into a few-step, frame-wise causal pipeline. Unlike [49, 13], which relies on heavy action blocks, we inject actions directly into the timestep embeddings, yielding a simpler and effective conditioning mechanism. At inference time, we maintain a context cache and generate frames autoregressively conditioned on both the cached context and incoming actions, enabling continuous, low-latency streaming generation. We evaluate under two settings: 1) With context memory: a window of $w$ frames is cached as clean world context in working memory, conditioning subsequent frame generation. 2) Without context memory: generation cold-starts from the initial image and proceeds autoregressively.
Table 2: Model Performance for the First Person on MIND-First 50. All videos underwent identical processing and were evaluated at 720p resolution. $\downarrow$ indicates lower values are better; $\uparrow$ indicates higher values are better.
[…]
w/o Context Memory (Image-to-World)
MIND-World (Ours) 0.1091 0.0359 0.1200 0.4583 0.5655 0.0356 0.4395
Matrix-Game 2.0 [13] 0.1188 0.0306 0.1084 0.4302 0.5180 0.0265 0.6914
w Context Memory (Video-to-World)
MIND-World (Ours) 0.1035 0.0309 0.1226 0.4590 0.5702 0.0384 0.5534
Table 3: Model Performance for the Third Person on MIND-Third 50. All videos underwent identical processing and were evaluated at 720p resolution. $\downarrow$ indicates lower values are better; $\uparrow$ indicates higher values are better.
[…]
w/o Context Memory (Image-to-World)
MIND-World (Ours) 0.1066 0.0327 0.0677 0.5204 0.5672 0.0271 0.2587
Matrix-Game 2.0 [13] 0.1404 0.0372 0.0777 0.4236 0.4857 0.0622 0.9031
w Context Memory (Video-to-World)
MIND-World (Ours) 0.1042 0.0316 0.0685 0.5300 0.5673 0.0321 0.3338
4.2 Experiment Setting
MIND-World. We initialize the foundation model with SkyReels-V2-I2V-1.3B [6] and fine-tune it with action injection for 33K steps with batch size $8$. For distillation, we initialize the $4$-step causal student based 1K teacher’s ODE trajectories and train for 33K steps, followed by 22K steps via DMD-based Self-Forcing. The student is strictly per-frame causal (chunk size = $1$) with a local attention window of 25 frames. All experiments are conducted on 44 $\times$ NVIDIA H100 GPUs.
MIND Dataset. As illustrated in Figure 5, the dataset covers $8$ major categories, comprising a total of $100$ first-person and $100$ third-person videos in the same action space, along with $25$ first-person and $25$ third-person videos in different action spaces. All videos are long-term, open-domain, high-quality, and frame-aligned. As illustrated in Figure 2, during the finetune training and test split, action distribution consistency is ensured, and the distribution across scene categories is balanced as much as possible. Among them, $50$ first-person and $50$ third-person videos from the shared action space are used for training MIND-World, while the remaining $150$ videos are reserved for the MIND evaluation.
4.3 Per-Dimension Evaluation
For each dimension, we compute the MIND score using the evaluation suite described in Section 3.5, with results presented in Tables 2 and 3. To advance future research, we introduce MIND-World as an open-domain video-to-world baseline with memory-augmented world modeling capabilities.
Memory Consistency. Tables 2 and 3 show that, on the long context memory metric, models with context memory outperform those without by more than $4\%$. generated scene consistency results further confirm the benefits of memory. Additionally, Matrix-game-2.0 performs poorly in third-person generation; human evaluation verifies that its metrics accurately reflect this limitation—the model fails to generate controllable third-person characters.
Action Accuracy. As shown in Tables 2 and 3, even when inputting context memory with the same action space as the fine-tuning phase, the world model’s action control performance still deteriorates. This indicates limitations in the current action injection mechanism, and how to design more effective action injection strategies to enhance the world model’s action control capability remains an important research problem worthy of in-depth exploration.
Action Space Generalization. Tables 2 and 3 show that injecting context memory impairs world model inference, since inconsistent action spaces disrupt reasoning in models without action generalization.
Visual Quality. Tables 2 and 3 collectively show that world models with memory produce videos of superior visual quality and better alignment with human aesthetic preferences. This is due to the memory mechanism leveraging richer contextual information, ensuring high consistency in style and coherence with the given segment.
Figure 7: Summary of insights from the challenges in MIND. For each challenge, a representative example is visualized.
4.4 Insights and Discussions
This section details the observations and insights derived from our comprehensive evaluation experiments.
Open Domain. As illustrated in Challenge $1$ of Figure 7, MIND-World trained on easily collected Minecraft datasets struggle to generalize to open-domain inference, whereas those trained on the high-quality dataset provided by MIND exhibit significantly improved generalization. However, acquiring such data is challenging; thus, effectively leveraging readily available large-scale data to achieve open-domain generalization remains a key open problem.
Action-Space Generalization. Tables 2 and 3 reveal that, in action space generalization, video-to-world models with memory capabilities underperform image-to-world models without memory. As shown in Challenge $2$ of Figure 7, further analysis indicates that memory-enabled models outperform memory-less ones within the original action space; however, their performance drops significantly when the action space changes. This suggests that context memory tied to an action space inconsistent with training disrupts model inference. Therefore, accurately detecting the action space from context memory and achieving precise prediction remains a major challenge.
Precise Action Control. In the experiment of Path $5$ in Figure 4, where the agent first moves left and then right to return to the starting position—the generated result is shown in Challenge 3 of Figure 7, Matrix-game-2.0 [13] fails to move left as expected, instead remaining stationary and ultimately stopping far to the right of the origin. In contrast, MIND-World correctly moves left but fails to return to the initial position after moving right. Repeated experiments reveal that the visual prompt (i.e., the input image or video) significantly affects action following. Therefore, separating visual prompts from action dynamics is key to achieving precise action control in world models.
Long-horizon Memory Consistency. Tables 2 and 3 show that, in long-horizon rollouts, models with memory significantly outperform those without. The visualization in Challenge $4$ of Figure 7 further confirms this: memory-enabled generations remain largely consistent with ground truth, while memory-less outputs deviate substantially. Moreover, current world models can only capture short-term memory; effectively maintaining and leveraging long-context memory remains a critical open problem.
Generated Scene Consistency. By conducting a mirroring experiment on Matrix-game-2.0 [13] along Path $5$ in Figure 4, the results as shown in Challenge $5$ of Figure 7 reveal that when the camera revisits previously generated scenes, the content is clearly inconsistent with prior generations. Therefore, ensuring consistent generation of scenes continues to pose a significant difficulty.
Third-person Perspective. As shown in Challenge $6$ of Figure 7, Matrix-game-2.0 [13] fails to control the third-person character and execute movement, causing the generated video to pass through the character and eventually lose it. In contrast, MIND-World can control the character but fails to properly handle the relationship between the foreground character and the background, resulting in the character passing directly through buildings. Therefore, accurately perceiving and modeling the interactions between characters and backgrounds remains a major challenge in world models.
5 Conclusion
We introduced MIND, the first open-domain closed-loop revisited benchmark for evaluating memory consistency and action control in world models from both first-person and third-person perspectives. Built on Unreal Engine 5 with diverse action spaces, MIND enables systematic assessment of long-term scene memory, temporal coherence, and action space generalization. Experiments with MIND-World reveal that the challenges remain in generalizing across action spaces and maintaining long-horizon coherence. MIND establishes a unified foundation for advancing interactive, temporally consistent open domain world model.